Circular No. 9113

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2010 B1 (CARDINAL) As first announced on CBET 2141, an apparently asteroidal object

discovered by R. D. Cardinal (Rothney Astrophysical Observatory, University of Calgary, 0.50-m f/1.0 reflector; discovery observation tabulated below), and placed on the ‘NEO Confirmation Page’, has been reported as cometary by several CCD astrometrists. Y. Ikari (Moriyama, Shiga-Ken, Japan, 0.26-m reflector) notes that some of his exposures on Jan. 22.6 UT showed a coma of diameter 0′.2. D. Chestnov, Moscow, writes that three stacked 300-s unfiltered exposures on Jan. 22.8 by T. Kryachko (Engelhardt Observatory, Zelenchukskaya Station, 0.3-m reflector) showed that the object appeared “softer” than the surrounding stars, having a 0′.1–0′.2 coma with a strong condensation but no tail. J. M. Aymami (Tiana, Spain, 0.25m reflector) remarks that 39 co-added images totaling 3900 seconds, taken around Jan. 22.9, showed a compact-but-nebulous object displaying what seemed to be a faint coma measuring 7′′.5.

2010 UT α_{2000} δ_{2000} Mag. Observer
Jan. 19.22247 8^{h}12^{m}28.^{s}90 +73^{o}42′01′′.1 17.7 Cardinal

The following parabolic orbital elements from MPEC 2010-B54 are from observations spanning Jan. 19–26:

T = 2011 Feb. 6.807 TT  ω = 211.^{o}674
Ω = 277.262 }2000.0
q = 2.93361 AU  i = 101.937

COMETS C/2009 M6, C/2009 M7, C/2009 M8, C/2009 N1 (SOHO)

Additional presumed comets have been found on SOHO website images — Kreutz sungrazers except for C/2009 M8 and C/2009 N1 (non-group). C/2009 M6 was very faint (mag ∼ 8.5) and slightly diffuse. C/2009 M7 was diffuse (mag ∼ 8). C/2008 M8 was appeared stellar, appearing quite bright (mag ∼ 6.5–7) but fading fast in C2 images. C/2009 N1 was slightly diffuse (mag ∼ 7) in C3 images, and a faint, diffuse streak in C2 images.

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 M6 June 25.604 6^{h}09.^{m}9 +21^{o}52′ C2  MU 2009-P02
C/2009 M7 26.908 6 15.3 +21 48 C2 BZ 2009-P20
C/2009 M8  30.071 6 32.4 +25 58 C3/2 RK 2009-P20
C/2009 N1 July 2.596 6 42.8 +26 41 C3/2 BZ 2009-P20
2010 January 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green